OpenStudy (anonymous):

Solve the triangle. B = 73°, b = 15, c = 10 @phi

OpenStudy (phi):

can you write down the law of sines for this problem ?

OpenStudy (anonymous):

\[\frac{ a }{ \sin A }=\frac{ b }{ \sin B }=\frac{ c }{ \sin C }\]

OpenStudy (phi):

yes, but we only 2 out of the 3. Leave out the a's what do we have ?

OpenStudy (phi):

*need 2 out of 3

OpenStudy (anonymous):

\[\frac{ b }{ \sin B }=\frac{ c }{ \sin C }\]?

OpenStudy (phi):

yes, now replace b, B and c with the numbers they gave you

OpenStudy (anonymous):

\[\frac{ 15 }{ \sin 73 }=\frac{ 10 }{ \sin C }\]

OpenStudy (phi):

now flip the fractions because we want sin C on top \[ \frac{ \sin 73 }{ 15 }=\frac{ \sin C }{ 10 } \] multiply both sides by 10. what do you get ?

OpenStudy (phi):

like this \[ \frac{ 10 \cdot \sin 73 }{ 15 }= \sin C \]

OpenStudy (anonymous):

\[\frac{ 9.563 }{ 150 }= \sin C\] I think I did it wrong. lol

OpenStudy (phi):

ok, here is how, in steps \[ \frac{10}{1} \cdot \frac{ \sin 73 }{ 15 }=\frac{ \cancel{10} \cdot \sin C }{ \cancel{10} }\]

OpenStudy (anonymous):

Would it be \[\frac{ 10\sin73 }{ 15 }=\sin C\]or\[\frac{ 9.563 }{ 15 }=\sin C\]

OpenStudy (phi):

you write down 10* (* for times) on both sides. 10 is 10/1 (to make it a "fraction") then multiply top * top to get 10 sin73 and bottom times bottom 1*15 = 15

OpenStudy (phi):

looks good, but divide by 15

OpenStudy (phi):

It's both of those ways 10 sin73/15 and 9.563/15 or even better 0.6375

OpenStudy (phi):

we want a number, so we can do the inverse sin to find angle C

OpenStudy (anonymous):

Would it be 39.61?

OpenStudy (phi):

if you mean angle C for "it" yes. we now have B= 73º, C= 39.61º all 3 angles in the triangle add up to 180º, so we can figure out angle A. what is it?

OpenStudy (anonymous):

A=70.39?

OpenStudy (phi):

I would add up B and C to get 73+39.61= 112.61 then 180 minus that for the third angle the check is A+B+C= 180

OpenStudy (anonymous):

Oh okay, so A=67.39 67.39+73+39.61=180

OpenStudy (phi):

Here is how to do it using algebra A+B+C= 180 replace B and C with their values A + 73+39.61 = 180 simplify the left side A + 112.61= 180 add -112.61 to both sides (or subtract 112.61 from both sides) A + 112.61-112.61= 180-112.61 A= 67.39

OpenStudy (phi):

what is left is to find side a. Use the Law of Sines again.

OpenStudy (anonymous):

\[\frac{ a }{ \sin 67.4 }=\frac{ 15 }{ \sin 73 }\]\[\frac{ \sin 67.4 }{ a }=\frac{ \sin 73 }{ 15 }\] Do I times it by 15?

OpenStudy (phi):

I would not flip the fraction, because a is on top. so use the first equation multiply both sides by sin 67.4 (that means write down sin67.4 don't use a calculator) don't evaluate (yet), just write the equation.

OpenStudy (anonymous):

\[\frac{ a }{ \sin 67.4 }(\frac{ 1 }{ \sin 67.4 })=\frac{ 15 }{ \sin 73 }\left( \frac{ 1 }{ \sin 67.4 } \right)\]\[a=\frac{ 15 }{ \sin73*\sin67.4 }\] Like this?

OpenStudy (phi):

almost. but we are not doing magic. you want to multiply by sin 67.4 not 1/sin67.4 notice that \[ \frac{a}{\sin 67.4} \frac{1}{\sin67.4} \] the sin * sin does not cancel. you want do divide sin/sin because anything divided by itself is 1 try again.

OpenStudy (phi):

lost ?

OpenStudy (anonymous):

\[\frac{ a }{ \sin67.4 }*\sin67.4\]It'll cancel them and just give us a, right?

OpenStudy (phi):

\[ \frac{ a }{ \sin 67.4 }(\frac{ \sin 67.4 }{ 1 })=\frac{ 15 }{ \sin 73 }\left( \frac{ \sin 67.4 }{ 1 } \right)\]

OpenStudy (phi):

yes, that is why we are multiplying both sides by sin 67.4

OpenStudy (anonymous):

\[a=\frac{ 15\sin67.4 }{ \sin73 }\]?

OpenStudy (phi):

yes, now use a calculator to find a

OpenStudy (anonymous):

14.5?

OpenStudy (phi):

yes, that looks good

OpenStudy (anonymous):

thank you

OpenStudy (phi):

Just to be complete, after we found the 2nd angle (angle C) to be 39.6, we should check for the "ambiguous case" if there was another solution (there isn't here, but we will check anyway) the other value of for C would be 180-39.6= 140.4 But when we add angle B = 73 + new angle C 140 we get 213, which is bigger than 180 so this is not an ambiguous case. C cannot be 140 degrees. But in some problems the 2nd angle 180-C will be small enough...